# Quantum-thermodynamic enhancements in continuous thermal machines require energetic coherence

### v1

@@ V.2 Thermodynamic disadvantage caused by degenerate coherence
We first introduce the following basis change where the degenerate levels |$\ket{2a}$ and |$\ket{2b}$ are transformed into levels |$\ket{\alpha}$ and |$\ket{\beta}$:
@@
Notice that in the case of symmetric rates for the hot reservoir, $\gamma_{\mathrm{h}}^{\mathrm{a}}=\gamma_{\mathrm{h}}^{\mathrm{b}}$, this expression becomes the one presented in Ref. [69], but differs otherwise. By introducing this change, we successfully decouple the state |$\ket{\beta}$ from the state |$\ket{0}$, resulting in a machine with only one explicit collective transition, as illustrated in Fig. 4a. For symmetric rates, we recover the situation where |$\ket{\beta}$ is completely decoupled from the states |$\ket{0}$ and |$\ket{1}$, leading to a classical three-level system with local jumps at same temperatures than the original NIC machine.
@@
Figure 4: (a.) Schematic representation of the level transitions in the Noise-induced machine after the variable change. (b.) Graphical representation of the inequality (35) separating the regimes where we can define the classical equivalent for noise-induced coherence (blue zone) and where we cannot (grey zone).
@@
In the general case, following Sec. IV.2, the classical equivalent includes an extra stochastic transition between the degenerate levels reading:
@@
$\begin{split}&\gamma_{\alpha\beta}^{\mathrm{cl}}=\gamma_{\beta\alpha}^{\mathrm{cl}}=\dfrac{(\bar{n}_{\mathrm{h}}+1)\gamma_{\mathrm{h}}^{\alpha}\gamma_{\mathrm{h}}^{\beta}}{(\bar{n}_{\mathrm{c}}+1)\gamma_{\mathrm{c}}^{\alpha}+(\bar{n}_{\mathrm{h}}+1)(\gamma_{\mathrm{h}}^{\alpha}+\gamma_{\mathrm{h}}^{\beta})},\\ \end{split}$ (33)
and corrections to the rates of the collective transitions between the degenerate levels and the other ones as
$\begin{split}\gamma_{k1}^{\mathrm{cl}}&=(\gamma_{\mathrm{h}}^{k}-2\gamma_{\alpha\beta}^{\mathrm{cl}})(\bar{n}_{\mathrm{h}}+1),\\ \gamma_{1k}^{\mathrm{cl}}&=(\gamma_{\mathrm{h}}^{k}-2\gamma_{\alpha\beta}^{\mathrm{cl}})\bar{n}_{\mathrm{h}},\end{split}$ (34)
@@
for $k=\alpha,\beta$. By examining the above equations, we find that the corrections to the spontaneous emission rates can make these rates eventually negative, which would correspond to a nonphysical situation. Consequently, the construction of the classical equivalent for noise-induced coherence is limited to scenarios that ensure positive rates. For the NIC machine analyzed here, this happens when
$\gamma_{\mathrm{h}}^{\alpha}+\gamma_{\mathrm{h}}^{\beta}\geq 2\mathrm{max}[\gamma_{\mathrm{h}}^{\alpha},\gamma_{\mathrm{h}}^{\beta}]-\dfrac{\bar{n}_{\mathrm{c}}+1}{\bar{n}_{\mathrm{h}}+1}\gamma_{\mathrm{c}}^{\alpha},$ (35)
as illustrated in Fig. 4b.
@@
We compute the fluctuation ratio $\mathcal{R}$ in Eq. (31) for the NIC machine and its classical equivalent by taking again $J_{\mathrm{out}}=\dot{W}$ for $10^{6}$ different parameter configurations within the region where the classical equivalent machine can be defined. In Fig. 5 we show the distribution of $\mathcal{R}$ together with the corresponding distribution of the TUR ratio $\mathcal{Q}_{\mathrm{he}}$ values (see inset). Our results show that $\mathcal{R}\geq 1$ for all cases, similarly to the case of the autonomous refrigerator. In particular, we find that the quantum-coherent machine is at least as noisy (or more) than its classical equivalent. Indeed for some configurations the NIC machine reaches noise levels up to $30\%$ higher than its classical counterpart. By looking at the inset we observe that, as expected, the TUR is also not violated for all choices of parameters ($\mathcal{Q}_{\mathrm{he}}\geq 2$ always) in accordance to previous studies about TUR violations in similar models [68].
@@
Figure 5: Histogram of sampled values of $\mathcal{Q}$ (inset plots) and the ratio between the fluctuations of the system and its classical equivalent $R$ (outset plots). The values corresponds to an exploration of the following region in the parameters space of the system: $\beta_{\mathrm{w}}\rightarrow\infty$, $\beta_{\mathrm{c}}=1$, $\beta_{\mathrm{h}}/\beta_{\mathrm{c}}\in[0,1]$, $\epsilon_{2}=5$, $\epsilon_{1}\in[0.1,4.9]$, $\gamma_{\mathrm{h}}^{(\mathrm{a}/\mathrm{b})}\in[10^{-5},10^{-2}]$ and $\gamma_{\mathrm{c}}^{(\mathrm{a}/\mathrm{b})}=10^{-3}$.
@@
Therefore it becomes clear that the NIC machine is not able to overcome the classical equivalent machine in any of the relevant configurations. Our results hence imply that noise-induced coherence has not any beneficial impact in the thermodynamic performance of the NIC machine, whose operation in terms of power, efficiency and reliability can be mimicked (or surpassed) by a purely classical equivalent machine using the same set of resources. These conclusions are in contrast with previous claims in the literature regarding quantum enhancements in NIC models, see e.g. Refs. [16, 19, 17, 18], which either lacked a systematic comparison to classical equivalent models, or were judged on the basis of more limiting notions of classical equivalent (e.g. requiring all model parameters to be maintained equal).
VI Pareto optimization and practical advantages
Figure 6: Two dimensional multi-objective optimization of the different thermodynamic quantities (normalized by its maximum) for the Three-level amplifier. The value of the fixed parameters are: $\beta_{\mathrm{c}}=1$, $\beta_{\mathrm{h}}=0.1$, $\epsilon_{2}=5$, $\omega_{\mathrm{d}}\in[0.1,4.9]$, $\gamma_{\mathrm{h}/\mathrm{c}}\in[10^{-5},10^{-2}]$ and $g\in[10^{-5},10^{-3}]$.
@@ VI Pareto optimization and practical advantages
In the previous section we have seen that only for the case of energetic coherence, corresponding to the case of the three-level amplifier introduced in Sec. II.1, a quantum-thermodynamic advantage can be identified for different set of parameters in the model. The reduction of the output current fluctuations as compared to classical equivalent machines operating at same (average) output power and efficiency, provide indeed necessary conditions for a quantum-thermodynamic advantage. Still in order to proof a practical (and hence useful) advantage, an analysis of the regimes where such an improved reliability is obtained is needed. Here we show that the advantages identified before occur indeed along optimal performance, e.g. for maximum power and for maximum efficiency regimes.
@@
Since optimizing the machine performance involves more than one desired objective (maximize power, maximize efficiency, minimize fluctuations), multi-objective optimization techniques, also called Pareto optimization, are required. These techniques are well known in engineering [80] and have been applied to a wide range of problems, from network theory [34] to betting strategies [35] and phenotypic response [36]. They have also attracted some attention recently in the field of stochastic and quantum thermodynamics, where they have been applied to the optimisation of heat engine cycles in the slow-driving regime [37, 38, 39].
@@
In order to solve a multi-objective optimisation problem (let’s take for example a two-dimensional problem) we first define a utility function that depends on the quantities to be optimized (e.g. maximisation of power and minimisation of fluctuations) with the form
@@
We applied the above optimization technique to the three-level amplifier for numerically obtaining Pareto fronts to three different multi-objective optimization problems, each of which optimizing two of the three performance indicators of the machine, namely average power $\langle\dot{W}\rangle$, efficiency $\eta$ and power fluctuations $\mathrm{Var}[\dot{W}]$. In Fig. 6 we can see the shape of the Pareto front (red curves) for the three different optimisation problems with fixed environmental temperatures and varying $\omega_{\mathrm{d}}$, $\gamma_{\mathrm{h}}$, $\gamma_{\mathrm{c}}$ and $g$. There, each circle (independently of its color) is generated from a possible configuration (choice of parameters) of the machine. For the optimization problems involving maximizing power $\langle\dot{W}\rangle$ and minimizing fluctuations $\mathrm{Var}[\dot{W}]$ (Fig. 6a) or maximizing power $\langle\dot{W}\rangle$ and minimizing efficiency $\eta$ (Fig. 6b), the Pareto fronts include a whole branch of solutions indicating a tradeoff between the optimization objectives. These go from the zero fluctuations point to the maximum power one in the first case, and from the Carnot point to maximum power point in the second one. On the other hand, for the problem of minimizing fluctuations $\mathrm{Var}[\dot{W}]$ and maximizing efficiency $\eta$(Fig. 6c) the Pareto front collapses to a single point, or optimal solution, achieving zero fluctuations and Carnot efficiency (but at zero output power).
@@
The role of energetic coherence in the performance of the engine can be unveiled by coloring the circles representing each parameter configuration according to their value of the TUR ratio $\mathcal{Q}_{\mathrm{he}}$ (orange for $\mathcal{Q}_{\mathrm{he}}\geq 2$ and blue for TUR violations, $\mathcal{Q}_{\mathrm{he}}<2$). As can be appreciated in Fig. 6, in all cases the configurations that violate the TUR span the domain classically achievable. More importantly, in Figs. 6a and b, we observe TUR violations within the Pareto fronts for large power outputs, and including the maximum power point. This result implies that three level amplifier shows a quantum-thermodynamic advantage (as spotted by the violation of the TUR) in relevant regimes of operation, where the machine performs in an optimal way.
Similarly, we can define and numerically solve the multi-objective optimization problems for the three-qubit refrigerator and NIC machines to obtain their respective Pareto fronts (see Appendix Y). Once all multi-objective optimization problems have been solved for the three quantum thermal machine models, we can compare their optimal solutions with the optimal solutions of their respective classical equivalents. A quantitative comparison of the optimal solutions (Pareto front) in each case, can be performed by computing the difference between the maximum values of their utility functions, $\Omega-\Omega_{\mathrm{cl}}$, for a given $\lambda$. We focus on the optimisation of $\langle J_{\mathrm{out}}\rangle$ and $\mathrm{Var}[J_{\mathrm{out}}]$, since by construction, the difference between the quantum machine and its classical equivalent lies in the variance of the currents. In Fig. 7 we show the behaviour of $\Omega-\Omega_{\mathrm{cl}}$ as a function of $\lambda$ for the three thermal machine models (and fixed bath temperatures). In all three cases, for small values of $\lambda$ i.e. close to the zero fluctuations point, we find no difference between the quantum and classical Pareto fronts. However, as soon as $\lambda$ increases, we find three quite different behaviours: (1) for the three-level amplifier the optimal solution is always better i.e. it reaches greater values of the utility function (that are indeed not achievable in the classical equivalent); (2) for the quantum absorption refrigerator it turns out that there is no difference at all between classical and quantum optimal solutions, hence being these machines exactly equivalent at the level of power, efficiency and stability; (3) the NIC machine is always less optimal than its classical counterpart, specially for high values of $\lambda$ (high powers) where the NIC machine becomes much noisier than its classical equivalent. In any case, we observe that in the limit $\lambda=1$ (maximum power point), for which fluctuations are not taken into account, the optimal solutions of quantum and classical models become equivalent, as expected by construction of the classical equivalent.
Figure 7: Difference between the quantum and classical maximums of the utility function $\Omega$ in the multi-objective optimization of $\langle\dot{W}\rangle-\mathrm{Var}[\dot{W}]$ as function of $\lambda$ for fixed environment temperatures (outset plot) and as a function of the environment temperatures for a fixed $\lambda=0.8$. The value of the fixed parameters are: $\beta_{\mathrm{c}}=1$, $\epsilon_{2}=5$, $\omega_{\mathrm{d}}\in[0.1,4.9]$, $\gamma_{\mathrm{h}/\mathrm{c}}\in[10^{-5},10^{-2}]$ and $g\in[10^{-5},10^{-3}]$.
Finally, also explore the mismatch in the utility functions, $\Omega-\Omega_{\mathrm{cl}}$, when varying the temperature bias (see the inset plot in Fig. 7). Fixing the value of $\lambda$ we observe that the advantages of the three-level amplifier quantum Pareto front are higher for high temperature biases and as we make $\beta_{\mathrm{h}}/\beta_{\mathrm{c}}$ smaller and smaller they tend to disappear, spotting again a (quantum) non-equilibrium effect. On the other hand, the disadvantages in the NIC machine front do not display a monotonic behavior with the temperature bias and tend to disappear for a high temperature bias, where the quantum and classical models become equally optimal.
VII Conclusions

### v2

@@ VI.3 Noise-induced coherent machine
We first introduce the following basis change within the coherent subspace, where the degenerate levels |$\ket{2a}$ and |$\ket{2b}$ are transformed into levels |$\ket{\alpha}$ and |$\ket{\beta}$, with:
@@
Notice that in the case of symmetric rates for the cold reservoir, $\gamma_{\mathrm{c}}^{\mathrm{a}}=\gamma_{\mathrm{c}}^{\mathrm{b}}$, this expression becomes the one presented in Ref. [69], but differs otherwise. By introducing this change, we successfully decouple the state |$\ket{\beta}$ from the state |$\ket{0}$, resulting in a machine with only one explicit collective transition, as illustrated in Fig. 5 (see Appendix for more detailed information).
@@
Figure 5: (a) Schematic representation of the level transitions in the NIC machine after the change of variables. Here one of the double coherent transitions are replaced by a simple jump (|$\ket{0}\leftrightarrow\ket{\alpha}$). (b) Graphical representation of the inequality (40) for $\gamma^{\rm a}_{\rm c}=\gamma^{\rm b}_{\rm c}=:\gamma_{\rm c}$, separating the regimes where we can define the classical equivalent for noise-induced coherence (blue zone) and where we cannot (grey zone). In the symmetric case (blue diagonal) the classical equivalent can always be defined.
@@
Following Sec. IV.2, the classical equivalent of the NIC machine includes an extra stochastic transition between the degenerate levels reading:
$\begin{split}&\gamma_{\alpha\beta}^{\mathrm{cl}}=\gamma_{\beta\alpha}^{\mathrm{cl}}=\dfrac{(\bar{n}_{\mathrm{h}}+1)^{2}\gamma_{\mathrm{h}}^{\alpha}\gamma_{\mathrm{h}}^{\beta}}{(\bar{n}_{\mathrm{c}}+1)\gamma_{\mathrm{c}}^{\alpha}+(\bar{n}_{\mathrm{h}}+1)(\gamma_{\mathrm{h}}^{\alpha}+\gamma_{\mathrm{h}}^{\beta})},\\ \end{split}$ (38)
together with corrections to the rates of the collective transitions
@@
$\begin{split}\gamma_{k1}^{\mathrm{cl}}&=\left(\gamma_{\mathrm{h}}^{k}-\dfrac{2\gamma_{\alpha\beta}^{\mathrm{cl}}}{\bar{n}_{\mathrm{h}}+1}\right)(\bar{n}_{\mathrm{h}}+1),\\ \gamma_{1k}^{\mathrm{cl}}&=\left(\gamma_{\mathrm{h}}^{k}-\dfrac{2\gamma_{\alpha\beta}^{\mathrm{cl}}}{\bar{n}_{\mathrm{h}}+1}\right)\bar{n}_{\mathrm{h}},\end{split}$ (39)
@@
for $k=\alpha,\beta$. By examining the above equations, we find that the corrections to the spontaneous emission rates can make these rates eventually negative, which would correspond to a nonphysical situation. Consequently, the construction of the classical equivalent for noise-induced coherence is limited to scenarios that ensure positive rates in Eq. (39). For the NIC machine analyzed here, this happens when
$\begin{split}\gamma_{\mathrm{h}}^{\beta}=0;&\;\;\;\gamma_{\mathrm{h}}^{\alpha}>0,\\ \gamma_{\mathrm{h}}^{\beta}>0;\;\;\;\gamma_{\mathrm{h}}^{\alpha}>\gamma_{\mathrm{h}}^{\beta};&\;\;\;\dfrac{(\gamma_{\mathrm{h}}^{\alpha}-\gamma_{\mathrm{h}}^{\beta})}{\gamma_{\mathrm{c}}^{\alpha}}\leq\dfrac{\bar{n}_{\mathrm{c}}+1}{\bar{n}_{\mathrm{h}}+1},\\ \gamma_{\mathrm{h}}^{\beta}>0;\;\;\;\gamma_{\mathrm{h}}^{\alpha}\leq\gamma_{\mathrm{h}}^{\beta};&\;\;\;\dfrac{(\gamma_{\mathrm{h}}^{\beta}-\gamma_{\mathrm{h}}^{\alpha})}{\gamma_{\mathrm{c}}^{\alpha}}\leq\dfrac{\bar{n}_{\mathrm{c}}+1}{\bar{n}_{\mathrm{h}}+1},\\ \end{split}$ (40)
which, taking into account the variable change together with $\bar{n}_{\mathrm{h}}\geq\bar{n}_{\mathrm{c}}$, corresponds to the stimulated emission rate in the cold bath being greater than or equal to the difference in stimulated emission rates in the hot bath between transitions from levels |$\ket{\alpha}$ and |$\ket{\beta}$.
@@
We obtain analytically the fluctuation ratio $\mathcal{R}$ in Eq. (31) for the NIC machine and its classical equivalent, taking again $J_{\mathrm{out}}=\dot{W}$. By evaluating the expression for $10^{6}$ different parameter configurations within the region where the classical equivalent machine can be defined, we show in Fig. 6 the distribution of $\mathcal{R}$ together with the corresponding distribution of the TUR ratio $\mathcal{Q}_{\mathrm{he}}$ values (see inset). Our results show that $\mathcal{R}\geq 0$ for some configurations, pointing to quantum-thermodynamic advantages in a similar way than the two previous examples. However, here we also get configurations leading to $\mathcal{R}<0$. For that configurations the quantum-coherent machine becomes more noisy than its classical equivalent, hence leading to a detrimental role for the coherence. Indeed for some configurations the NIC machine reaches noise levels up to $1\%$ higher than its classical counterpart. Moreover, by looking at the inset we observe that, as expected, the TUR is not violated for all choices of parameters ($\mathcal{Q}_{\mathrm{he}}\geq 2$ always) in accordance to previous studies about TUR violations in similar models [68].
@@
Figure 6: Histogram of sampled values of $\mathcal{Q}$ (inset plots) and the ratio between the fluctuations of the system and its classical equivalent $R$ (outset plots). The values corresponds to an exploration of the following region in the parameters space of the system: $\beta_{\mathrm{w}}\rightarrow 0$, $\beta_{\mathrm{c}}=1$, $\beta_{\mathrm{h}}/\beta_{\mathrm{c}}\in[0,1]$, $\epsilon_{2}=5$, $\epsilon_{1}\in[0.1,4.9]$, $\gamma_{\mathrm{h}}^{(\mathrm{a}/\mathrm{b})}\in[10^{-5},10^{-2}]$ and $\gamma_{\mathrm{c}}^{(\mathrm{a}/\mathrm{b})}=10^{-3}$ (energies are given in $k_{\rm B}T_{\rm c}=1$ units).
Figure 7: Violations of the TUR and two dimensional multi-objective optimization for power $\langle\dot{W}\rangle$, power fluctuations ${\rm Var}[\dot{W}]$ and efficiency $\eta$ for the Coherent three-level amplifier. Points are generated from $10^{5}$ different configurations in the parameter space $\omega_{\mathrm{d}}\in[0.1,4.9]$, $\gamma_{\mathrm{h}/\mathrm{c}}\in[10^{-5},10^{-2}]$ and $g\in[10^{-5},10^{-3}]$ and fixed $\beta_{\mathrm{c}}=1$, $\beta_{\mathrm{h}}=0.1$, $\epsilon_{2}=5$ (energies are given in $k_{\rm B}T_{\rm c}=1$ units). Blue points represent configurations that violate the classical TUR and Pareto fronts to the red dashed line (see legend). Axes are normalized by the respective maximum values $\langle\dot{W}\rangle_{\rm max}=9.5\times 10^{-4}k_{B}T_{\rm c}/\hbar$, ${\rm Var}[\dot{W}]_{\rm max}=5.1\times 10^{-3}(k_{B}T_{\rm c}/\hbar)^{2}$, and $\eta_{C}=0.9$.
@@
Therefore, it becomes clear that the NIC machine is able to overcome the classical equivalent machine in some configurations which are not witnessed by TUR violations as in the case of Hamiltonian-induced coherence between degenerate levels; but not in all configurations. Contrary to the case of Hamiltonian-induced coherence, noise-induced one has not necessarily a beneficial impact in the thermodynamic performance of the NIC machine, whose operation in terms of power, efficiency and reliability can be surpassed in some regimes by a purely classical equivalent machine using the same set of resources.
The region of parameters in which a quantum-thermodynamic advantage over the classical equivalent is verified can be obtained analytically (see App. C) and is given by:
$\dfrac{\gamma_{\mathrm{h}}^{\alpha}}{\gamma_{\mathrm{h}}^{\beta}}\geq\dfrac{2+3(\bar{n}_{\mathrm{h}}+\bar{n}_{\mathrm{c}})+4\bar{n}_{\mathrm{h}}\bar{n}\mathrm{c}}{\bar{n}_{\mathrm{h}}-\bar{n}_{\mathrm{c}}},\;\;\;\gamma_{\mathrm{h}}^{\beta}\neq 0.$ (41)
Detrimental effects from coherence are obtained for regions where the above inequality is inverted. In particular, in the symmetric case $\gamma_{\rm c}^{\rm a}=\gamma_{\rm c}^{\rm b}$ and $\gamma_{\rm h}^{\rm a}=\gamma_{\rm h}^{\rm b}$ we will have that $\gamma_{\mathrm{h}}^{\beta}=0$ and the system effectively becomes a three level system without coherence in the steady state. In this limit, the quantum system and its classical analog become indistinguishable. These conclusions are in contrast with some previous claims in the literature claiming quantum enhancements (also in the symmetric case) for same or similar NIC models, see e.g. Refs. [16, 19, 17, 18], which lacked a careful comparison to classical thermodynamic-equivalent models.
VII Pareto optimization and practical advantages
@@
In the previous sections we have seen that only for the case of Hamiltonian-induced coherence, corresponding e.g. to the cases of the three-level amplifier introduced in Sec. II.1 and to the three-qubit refrigerator in Sec. II.2, a quantum-thermodynamic advantage is guaranteed in out-of-equilibrium conditions for all range of parameters. In that case, there is a net reduction of the output current fluctuations as compared to classical equivalent machines operating at same (average) output power and efficiency, hence providing performances not achievable with classical stochastic machines using the same amount of resources. Still in order to proof a practical (and hence useful) advantage, an analysis of the regimes where such an improved reliability is obtained is needed. Here we show that the advantages identified before occur indeed along optimal performance, e.g. for maximum power and for maximum efficiency regimes.
@@
Since optimizing the machine performance involves more than one desired objective (maximize power, maximize efficiency, minimize fluctuations), multi-objective optimization techniques, also called Pareto optimization, are required. These techniques are well known in engineering [93] and have been applied to a wide range of problems, from network theory [34] to betting strategies [35] and phenotypic response [36]. They have also attracted some attention recently in the field of stochastic and quantum thermodynamics, where they have been applied to the optimization of heat engine cycles in the slow-driving regime [37, 38, 39].
@@
In order to solve a multi-objective optimization problem (let’s take for example a two-dimensional problem) we first define a utility function that depends on the quantities to be optimized (e.g. maximization of power and minimization of fluctuations) with the form
@@
We applied the above optimization technique to the prototypical models introduced in Sec. II for numerically obtaining Pareto fronts to three different multi-objective optimization problems, each of which optimizing two of the three performance indicators of the machine, namely average output power $\langle\dot{W}\rangle$, efficiency $\eta$ and power fluctuations $\mathrm{Var}[\dot{W}]$. For illustrative purposes we focus first on the three-level amplifier of Sec. II.1, and then on the three-qubit autonomous refrigerator of Sec. II.2.
@@
In Fig. 7 we can see the shape of the Pareto front (red curves) of the three-level amplifier for the three different optimization problems with fixed environmental temperatures and varying $\omega_{\mathrm{d}}$, $\gamma_{\mathrm{h}}$, $\gamma_{\mathrm{c}}$ and $g$. There, each circle (independently of its color) is generated from a possible configuration (choice of parameters) of the machine. For the optimization problems involving maximizing power $\langle\dot{W}\rangle$ and minimizing fluctuations $\mathrm{Var}[\dot{W}]$ (Fig. 7a) or maximizing power $\langle\dot{W}\rangle$ and minimizing efficiency $\eta$ (Fig. 7b), the Pareto fronts include a whole branch of solutions indicating a trade-off between the optimization objectives. These go from the zero fluctuations point to the maximum power one in the first case, and from the Carnot point to maximum power point in the second one. On the other hand, for the problem of minimizing fluctuations $\mathrm{Var}[\dot{W}]$ and maximizing efficiency $\eta$ (Fig. 7c) the Pareto front collapses to a single point, or optimal solution, achieving zero fluctuations and Carnot efficiency (but at zero output power).
Figure 8: Fluctuations ratio and two dimensional multi-objective optimization of power $\langle\dot{W}\rangle$, power fluctuations ${\rm Var}[\dot{W}]$ and efficiency $\eta$ for the Coherent three-level amplifier. Points are generated from $10^{5}$ different configurations in the parameter space $\omega_{\mathrm{d}}\in[0.1,4.9]$, $\gamma_{\mathrm{h}/\mathrm{c}}\in[10^{-5},10^{-2}]$ and $g\in[10^{-5},10^{-3}]$ and fixed $\beta_{\mathrm{c}}=1$, $\beta_{\mathrm{h}}=0.1$, $\epsilon_{2}=5$ (energies are given in $k_{\rm B}T_{\rm c}=1$ units). The color scale stands for the fluctuation ratio $\mathcal{R}$ and Pareto fronts are highlighted in red (see legend). Axes are normalized by the respective maximum values $\langle\dot{W}\rangle_{\rm max}=9.5\times 10^{-4}k_{B}T_{\rm c}/\hbar$, ${\rm Var}[\dot{W}]_{\rm max}=5.1\times 10^{-3}(k_{B}T_{\rm c}/\hbar)^{2}$, and $\eta_{C}=0.9$.
@@
The role of coherence in the performance of the engine can be unveiled by coloring the circles representing each parameter configuration according to their value of the TUR ratio $\mathcal{Q}_{\mathrm{he}}$ (orange for $\mathcal{Q}_{\mathrm{he}}\geq 2$ and blue for TUR violations, $\mathcal{Q}_{\mathrm{he}}<2$). As can be appreciated in Fig. 7, in all cases the configurations that violate the TUR span the domain classically achievable. More importantly, in Figs. 7a and 7b, we observe TUR violations within the Pareto fronts for large power outputs, and including the maximum power point. This result implies that the coherent three-level amplifier shows a quantum-thermodynamic advantage (as spotted by the violation of the TUR) in relevant regimes of operation, where the machine performs in an optimal way. The TUR violations in Fig. 7 are to be compared with the value of the fluctuations ratio $\mathcal{R}$ for the same parameter configurations in the optimization problems, as shown in Fig. 8 for power and fluctuations (a), and power and efficiency (b). As predicted by Theorem 1 in Sec. V, we obtain positive values $\mathcal{R}>0$ except at zero power, when equilibrium is reached. Moreover, the presence of large values of $\mathcal{R}$ along the Pareto front confirm the presence of practical quantum-thermodynamic advantages in optimal regimes of operation for $\langle\dot{W}\rangle>0$, where the largest values of $\mathcal{R}$ are obtained at the maximum power point.
Figure 9: Fluctuations ratio and two dimensional multi-objective optimization of cooling power $\langle\dot{Q}_{\rm c}\rangle$, its fluctuations ${\rm Var}[\dot{Q}_{\rm c}]$ and the coeficient of performance $\eta$ for the three-qubit autonomous refrigerator. Points are generated from $10^{5}$ different configurations in the parameter space $\epsilon_{1}\in[0.1,4.9]$, $\gamma_{\mathrm{h}/\mathrm{m}/\mathrm{c}}\in[10^{-5},10^{-2}]$ and $g\in[10^{-5},10^{-3}]$ and fixed parameters $\beta_{\mathrm{c}}=1$, $\beta_{\mathrm{m}}/\beta_{\mathrm{c}}=0.5$, $\beta_{\mathrm{h}}/\beta_{\mathrm{c}}=0.05$, $\epsilon_{2}=5$ (energies are given in $k_{\rm B}T_{\rm c}=1$ units). Axes are normalized by the respective maximum values $\langle\dot{Q}_{\rm c}\rangle_{\rm max}=2.9\times 10^{-4}k_{B}T_{\rm c}/\hbar$, ${\rm Var}[\dot{Q}_{\rm c}]_{\rm max}=2.4\times 10^{-3}(k_{B}T_{\rm c}/\hbar)^{2}$, and $\eta_{\rm abs}=0.9$
On the other hand, for the three-qubit autonomous refrigerator, the TUR is never violated and hence it cannot be used to spot practical advantages in optimal operation regimes. In this case, our method is nevertheless able to show that practical quantum-thermodynamic advantages can be obtained despite the absence of TUR violations, as predicted by Theorem 1 in Sec. V. In Fig. 9 we show the colormap of $\mathcal{R}$ for the corresponding two optimization problems of the three-qubit refrigerator with fixed environmental temperatures and varying $\omega_{\mathrm{d}}$, $\gamma_{\mathrm{h}}$, $\gamma_{\mathrm{c}}$ and $g$. The corresponding Pareto fronts are shown again in red color, which have qualitatively similar shapes as in the coherent three-level amplifier [Fig. 8]. Also for this case, Fig. 9 reveals that configurations with values of $\mathcal{R}>0$ are obtained along the Pareto fronts, confirming the presence of quantum-thermodynamic advantages produced by the presence of degenerate coherence in optimal regimes of operation. In this case, maximum values appear along all the Pareto front ${\rm Var[\dot{Q}_{\rm c}]}-\langle Q_{\rm c}\rangle$. However, we observe that the values of $\mathcal{R}$ are in general much smaller than in the previous case, showing a higher impact of energetic coherence to produce quantum advantages in comparison to degenerate one.
VIII Conclusions